One pound of apples costs $2 and one pound of peaches costs $1. How many pounds of each fruit do I need to buy so that I have 20 pounds of fruit and spend exactly $28? (I've tried everything I can think of!!)
Answer by hamsanash1981@gmail.com(151)

Let pounds of apple be x and peaches be y,
cost of 1 pound of apple = 2 and cost of peach is 1
then the cost of X pounds of apple = 2x and cost of Y pounds of peach = y
total pounds is 20 => x + Y = 20 --- 1
and total cost is 28 => 2x + y = 28 --2
Solving equ 1 and 2
x + Y =20
2x + y = 28
===========
x = 8
then Y = 12
Therefore 8 pounds of apple and 12 pounds of peach is needed to buy to spend exactly $28.